An earth satellite has a mass M and angular momentum L.Its areal velocity is given by........

Grade:Select Grade

1 Answers

Saurabh Kumar
9 years ago
Areal velocity, i.e. dA/dt,i.e. d(1/2Arc x Angle)/dt = d(1/2R2angle)/dt.
Since, the radius is constant but the differentiation of angle is angular velocity.
1/2R2w = L/2MR2w